15 October 2013

15th of October

Farm news
We had a few hot days again this week where the farm was completely off.

Late last week I did a UPS test by switching off the power board that its plugged into to see if the server continued running. It did. After 10 minutes I turned it back on to find the KVM was no longer working. After a number of attempts to get it going I have had to take it back to the shop. New ones cost $600 and the distributor doesn't have stock until next month. I have asked to get it repaired and of course its not under warranty. Because the other cluster of machines are off I have swapped the other KVM over to get going.

The Intel-GPU machines are still finishing off their climate models. The Raspberry Pi's are running Einstein work because Asteroid ran out of work.

BOINC testing
We got up to 7.2.21 early this week (or .22 on the Mac). The main change with this is OpenCL support of CPUs.

Project news - Climate Prediction
I still can't upload. This week their server is reporting its out of disk space. As well as that I can't report any completed work.

I have ordered a 56k dial-up modem which I will use via another ISP in case of issues like this.

No comments: